Two long, straight, parallel wires 24 cm apart carry currents of I1 = 30A andI2 = 20A in the same direction. Determine the magnetic field in the plane of the twowires at a point halfway between the wires.
1
Expert's answer
2019-03-20T11:47:01-0400

Induction of the direct wire field B = μ0I/2πr, in accordance with the principle of superposition Bh = μ0 (I1 - I2)/πl = 4*10-7*10/ 0.24 = 16.7 (mT).

So, B = 16.7 mT.